# Price Experimentation and Interference

### v3

@@
In this paper, we examine biases arising in A/B tests where firms modify a continuous parameter, such as price, to estimate the global treatment effect of a given performance metric, such as profit. These biases emerge in canonical experimental estimators due to interference among market participants. We employ structural modeling and differential calculus to derive intuitive characterizations of these biases. We then specialize our general model to a standard revenue management pricing problem. This setting highlights a key pitfall in the use of A/B pricing experiments to guide profit maximization: notably, the canonical estimator for the expected change in profits can have the wrong sign. In other words, following the guidance of canonical estimators may lead firms to move prices in the wrong direction, inadvertently decreasing profits relative to the status quo. We introduce a novel debiasing technique for such experiments, that only requires that the firm equally split experimental units between treatment and control. We apply these results to a two-sided market model and show how this “change of sign” regime depends on model parameters such as market imbalance, as well as the price markup. We also extend our revenue management model to an alternative specification in which firms set a fixed sales fee across heterogeneous products, and show that all our results extend to this setting as well.
[…]
@@ 1 Introduction
In this work we analyze the biases that arise when a technological firm or platform conducts A/B tests by varying a continuous parameter to determine the global treatment effect associated to a performance metric of interest. These biases appear when using canonical designs and estimators and result from interference between market participants.
@@
More concretely, consider the example of a canonical pricing experiment where a fraction of the products are assigned a price that is higher than the baseline price. The global treatment effect of this higher price compares two counterfactual universes: one where all products have the high price (global treatment), and one where all products have the low price (global control). Observe that in the experiment, some consumers who would have purchased at the high price in global treatment instead purchase control units at the low price; when using canonical estimators (such as the difference-in-means estimator), this substitution within the experiment biases the estimation of the global treatment effect. This phenomenon is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts; see Related Work (Section 2) for background.
@@
Our paper considers the following question: Given the standard experimental designs (e.g., product-level or customer-level randomization) and estimators (i.e., difference-in-means) that are employed by platforms, what are the consequences of interference bias in pricing experiments? There are several key features of our approach. First, drawing inspiration from recent work in the operations literature, we employ structural modeling to gain valuable insights into biases. Furthermore, our analysis focuses on differential changes in continuous variables such as prices, enabling us to leverage calculus to derive straightforward and intuitive expressions. (In practical settings, we will think of “small” changes around the control value.) Finally, unlike many preceding studies on online platform experimentation, we focus on profits as an outcome metric, in addition to bookings or transaction volume.
@@
We first introduce a succinct yet general model of platform experimentation in which a firm wants to understand the effect of a continuous variable on a metric of interest. The model is general in that can it encode different market arrangements and experimental designs. By leveraging differential calculus, we provide several structural results regarding the performance of canonical estimators and designs. Notably, we show that the bias corresponds to the cross-effects of the control and treatment metrics with respect to marginal changes on the other group’s variable. These cross-effects are precisely an indicator of interference, providing an intuitive structural characterization of bias.
@@
Second, we specialize our results to a canonical revenue management pricing problem. By using our bias characterizations, we show that experiments can obtain the wrong sign when the metric of interest is profit (or revenue). In particular, we show that when treatment involves raising prices (resp., lowering prices), the global treatment effect may be positive (resp., negative), but due to interference, the estimator can be negative (resp., positive). (Interestingly, when treatment involves raising prices (resp., lowering prices), and the global treatment effect is negative (resp., positive), the experimental estimate obtains the correct sign.) We express the conditions under which an estimate with incorrect sign can be obtained with straightforward expressions that involve price elasticities. [^1] The expressions are reminiscent of elasticity conditions in Blake and Coey, 2014a . Our results have important implications for price optimization based on experiments: incorrectly following the experiment’s suggestion may lead one to move in the wrong direction and decrease profits relative to the status-quo!
@@
We take advantage of our structural characterization of bias to construct a novel debiasing method for the canonical A/B experiment. In particular, we show that if the experiment equally allocates units between treatment and control, then it is possible to fully debias the canonical estimator, by taking advantage of symmetry in the structure of bias. Notably, this allows firms to run the most simple version of an A/B pricing experiment, while still emerging with an unbiased estimator (under a set of assumptions that me make clear below). This is a stark contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require much more complicated experimental designs and/or estimators (Holtz et al., 2020, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
@@
To obtain additional insight, we apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022). We show how the biases of listing and customer-side experiments depend on model parameters with a particular focus on regimes in which this “change of sign” occurs. For example, for a listing-side experiment, a platform is more likely to obtain the incorrect sign in an experiment if the demand is relatively low. We also investigate numerically how bias depends on both market balance and the baseline control price.
We conclude the paper with an alternative specialization of our revenue management setting to a model where the firm sets a fixed fee as a percentage of overall transaction volume, as is common in online marketplaces. In this model we allow the possibility of a heterogeneous collection of goods for sale, with an unknown (joint) demand function. We show that all our previous results extend to this setting as well.
@@ 2 Related Work
The phenomenon we study is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts. More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015). Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014b , Fradkin, 2019, Liu et al., 2020), showing its importance in practice. Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), and message passing (shirani2023causal).
@@
A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), roemheld2023interference, Deng and Le, (2023). Both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects). Roemheld and Rao, (2024) uses an empirical study to suggest false positives may be inflated due to interference in pricing experiments. More similar to our findings, Deng and Le, (2023) also notes the possibility that the sign of the treatment effect can flip; in their setting this is due to an effect they call lead-day bias in markets with expiring goods. In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign. Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper – those where a continuous variable (such as price) is incrementally varied – as marginal policy changes, and develops a framework to study them.
@@
The closest of these works to ours is Munro et al., (2021), which studies general estimands in an economic price equilibrium setting, and decomposes the treatment effect into direct and indirect effects; the former is obtained via “naive” estimators, while the latter corresponds to our bias, and the authors show how it can be corrected using local price experiments. In that work, the direct and indirect effects are obtained as marginal changes with respect to the treatment assignment probability (inspired by work in Hu et al., (2022)), rather than the price in our present paper. Further, our emphasis in applying our results is on structural insight regarding when changes of sign may be obtained for objectives such as profits or revenue, while their paper is more focused on design and estimation for general estimands in market equilibrium. (An earlier version of Munro et al., (2021) considers treatment changes in a continuous variable, with an emphasis on relating the resulting differential treatment effect to direct and indirect effects.)
@@ 3.1 Market Model and Estimand
The firm is interested in optimizing the performance metric by adjusting the decision variable $x$. In particular, the firm is interested in the difference in the metric when the decision variable takes on a new “treatment” value, denoted by $x_{1}$, relative to a status-quo “control” value, denoted by $x_{0}$. This difference, $T(x_{1})-T(x_{0})$, corresponds to the global treatment effect ($GTE$). We define a local version of $GTE$ considering that the treatment value of the decision variable is “close” to the control value:
$GTE(x_{0})$ ::=:= $T^{\prime}(x_{0})\ .$ (1)
@@ 3.2 Experiment: Design, Model, and Estimator
Our work focuses on settings where the function $T$ may be unknown to the firm or platform. Therefore, an experiment is conducted to determine the direction in which the decision variable should be modified to optimize the performance metric. We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction $1-q$ to control. We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated in each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
@@
It is worth noting that in the settings we are considering, randomization can occur at either side of the market, i.e., customers or products (i.e., listings in the case of a platform) can be randomized. In a customer-side experiment, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. On the other hand, in a product (listing)-side experiment, a customer is exposed to a mix of treatment and control products (or listings). As before, we denote by $x_{0}$ the baseline (control) value of the decision variable and by $x_{1}$ the treatment value.
@@
We assume that the performance metric under the treatment and control conditions are given by the functions $T_{0}(x_{0},x_{1},q)$ and $T_{1}(x_{0},x_{1},q)$, which we refer to as the control and treatment performance metrics, respectively. The function $T_{0}(x_{0},x_{1},q)$ (resp., $T_{1}(x_{0},x_{1},q)$) corresponds to the performance metric in the control (resp., treatment) group when the control and treatment value are given by $x_{0}$ and $x_{1}$, respectively, and a fraction $q$ of units is randomized into treatment. The nature of these functions will become more clear in the context of the applications described below. We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
@@
The canonical experimental designs at each side of the market have an associated natural “naive” local estimator. To define this estimator, consider a sequence of experiments for which the treatment value of the decision variable approaches the control value. Then, we compute the relative difference between the performance metrics appropriately normalized by the treatment and control fractions, respectively, which corresponds to the Horvitz-Thompson estimator, or equivalently in our formal mean-field setting, the difference-in-means estimator. More formally, we consider a sequence of experiments indexed by $n\in\mathbb{Z}_{+}$, for which the treatment approaches control: $\lim_{n\to\infty}x_{1}^{n}=x_{0}$. We define a local version of the naive estimator as:
$\widehat{GTE}(x_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$ (2)
@@
The assumption states that if the values of control and treatment decision variables are the same, then both the scaled control and treatment performance metrics correspond to the market-wide metric. The assumption will play a key role in our analysis. To further understand it, consider $T_{1}$ with an extended domain $0<q\leq 1$ (the argument for $T_{0}$ is symmetric). Now, suppose $T_{1}$ is linear in its third argument over the relevant domain, that is, for any $0<q\leq 1$ and $10<\alpha<$ 1, $T_{1}(x,x,\alpha q)=\alpha T_{1}(x,x,q)$. If we take $q=1$, then $T_{1}(x,x,\alpha)/\alpha=T_{1}(x,x,1)=T(x)$, where the second equation follows directly by definition. Hence, the assumption is implied by constant (i.e., linear) returns of the performance metric in the fraction randomized. Note that experimental estimates in practice typically implicitly rely on the assumption of constant returns to extrapolate estimates for control and treatment groups to the entire market. For the marketplace models and metrics we study later, this assumption will be satisfied.
@@
$\widehat{GTE}(x_{0},q)$ $=$ $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{1}(x_{0},x_{0},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$ (3)
@@
Finally, we define the estimation bias of this local estimator as:
$Bias(x_{0},q)$ ::=:= $GTE(x_{0})-\widehat{GTE}(x_{0},q)$ (4)
$=$ $T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)$ (5)
@@
The result provides an intuitive characterization of bias. Specifically, $T_{0y}(x_{0},x_{0},q)$ and $T_{1x}(x_{0},x_{0},q)$ correspond to the relative changes in performance for the control and treatment groups, respectively, when the decision variable on the other group changes marginally. The cross-partial derivatives represent the effect on the performance metric across experimental groups, making them an intuitive indicator of interference. As interference increases, bias correspondingly intensifies.

### v4

@@
In this paper, we examine the biases that arise when firms run A/B tests on continuous parameters to estimate global treatment effects on performance metrics of interest; we particularly focus on price experiments to measure the price impact on quantity demanded, and on profit. In canonical A/B experimental estimators, biases emerge due to interference between market participants. We employ structural modeling and differential calculus to derive intuitive characterizations of these biases. We then specialize our general model to the standard revenue-management pricing problem. This setting highlights a fundamental risk innate to A/B pricing experiments: that the canonical estimator for the expected change in profits, counterintuitively, can have the wrong sign in expectation. In other words, following the guidance of canonical estimators may lead firms to move prices (or fees) in the wrong direction, inadvertently decreasing profits. We introduce a novel debiasing technique for these canonical experiments, requiring only that firms equally split units between treatment and control. We apply these results to a two-sided market model, and demonstrate how the “change of sign” regime depends on market factors such as the supply/demand imbalance, and the price markup. We conclude by calibrating our two-sided market model to published empirical estimates from Airbnb marketplaces, demonstrating that estimators with the wrong sign are not a knife-edge issue, and that they may be prevalent enough to be of concern to practitioners. [^1] We would like to thank Wassim Dhaouadi for his valuable input during the beginning stages of the project, and Andrey Fradkin for his critical assistance for our Airbnb empirical case study. We are also grateful to Eric Aldrich, Alex Deng, Thu Le, Ruben Lobel, Ali Rauh, Dan Schmierer, and seminar participants at Boston University, University of Chile, the MIT Conference on Digital Experimentation, the INFORMS Revenue Management and Pricing Conferencea, the INFORMS Annual Meeting, the INFORMS Manufacturing and Service Operations Management Conference, and Airbnb for helpful feedback. Orrie Page was supported by a National Science Foundation Graduate Research Fellowship.
@@ 1 Introduction
In this work, we analyze the biases that arise when a firm or platform conducts an A/B test by varying a continuous parameter, in order to determine the global treatment effect associated with a performance metric of interest. These biases appear when using canonical experimental designs and estimators, and are the result of interference between market participants.
@@
More concretely, consider the example of a typical pricing experiment, commonly used by online platforms, where a fraction of the products are assigned a treatment price that is higher than the baseline (control) price. The global treatment effect of this higher price compares two counterfactual universes: one where all products have the high price (global treatment), and one where all products have the low price (global control). Observe that in the experiment, some consumers who would have purchased at the high price in global treatment instead purchase control units at the low price. When using canonical estimators (such as the difference-in-means estimator), this substitution within the experiment biases the estimation of the global treatment effect. This phenomenon is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts; see Related Work (Section 2) for background.
@@
Our paper considers the following question: Given the standard experimental designs (e.g., listing-side randomization, LR, or customer-side randomization, CR) and estimators (i.e., difference-in-means) that are employed by platforms, what are the consequences of interference bias in pricing experiments? There are several key features of our approach. First, drawing inspiration from recent work in the operations literature, we employ structural modeling to gain valuable insights into biases. Furthermore, our analysis focuses on differential changes in continuous variables such as prices, enabling us to leverage calculus to derive straightforward and intuitive expressions. (In practical settings, this corresponds to “small” changes around the control price.) Finally, unlike many preceding studies on online platform experimentation, we focus on profits as an outcome metric, in addition to bookings or transaction volume.
@@
We first introduce in Section 3 a succinct yet general model of platform experimentation in which a firm wants to understand the impact of modifications to a continuous variable, on a metric of interest. The model is general, in the sense that it can represent various market arrangements and experimental designs. By leveraging differential calculus, we provide several structural results regarding the performance of canonical estimators and designs. Notably, we show that the bias corresponds to the cross-effects of the control and treatment group metrics, with respect to marginal changes in the other group’s variable. These cross-effects are precisely an indicator of interference, providing an intuitive structural characterization of bias.
@@
Second, in Sections 4 and 5, we specialize our results to a canonical revenue management pricing problem. By using our bias characterizations, we show that the bias can be severe enough that experimenters can obtain the wrong sign when the metric of interest is profit (or revenue). In particular, we show that when treatment involves raising prices (resp., lowering prices), the global treatment effect may be positive (resp., negative), but due to interference, the estimator can be negative (resp., positive) in expectation. Interestingly, when treatment involves raising prices (resp., lowering prices), and the global treatment effect is negative (resp., positive), the experimental estimate always obtains the correct sign. We express the conditions under which an estimate with the wrong sign can be obtained with straightforward expressions that involve price elasticities. [^2] The expressions are reminiscent of elasticity conditions in Blake and Coey, (2014), though their analysis differs from ours: they focus on estimating the relative bias in terms of supply and demand elasticities, and carry out a local change with respect to a multiplicative factor by which treatment demand changes relative to control. By contrast, our estimand is local with respect to the magnitude of the price change, and of course, a significant emphasis of our study is the possibility of mismatched signs between the true $GTE$ and the estimated $GTE$, which they do not study. In the appendix, we also show that these results generalize to a setting where a platform with heterogeneous goods for sale sets a fixed fee as a percentage of overall transaction volume, as is common in online marketplaces. Our results have important implications for price optimization based on experiments: following the suggestion of a canonical experiment may lead one to move in the wrong direction and decrease profits relative to the status-quo!
@@
Taking advantage of our structural characterization of bias, in Section 6, we construct a novel debiasing method for canonical A/B experiments. In particular, we show that if the experimenter equally allocates units between treatment and control, it is possible to fully debias the canonical estimator by taking advantage of symmetry in the structure of bias. Notably, this allows firms to run the most simple version of an A/B pricing experiment, while still emerging with an unbiased estimator (under a set of assumptions that me make clear below). This is a contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require more complicated experimental designs and estimators (Holtz et al., 2020, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
@@
To obtain additional insight, in Section 7, we apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022). We show how the biases of LR and CR experiments depend on model parameters, with a particular focus on regimes in which a “change of sign” may occur. For example, for an LR experiment, a platform is more likely to obtain the incorrect sign in an experiment if the demand is low relative to supply. We also investigate numerically how bias depends on both market balance and the baseline control price, and show how canonical A/B methods for price experimentation may result in significant profit underperformance.
In Section 8, we conclude the paper by calibrating our two-sided market model to empirical estimates from Farronato and Fradkin, (2022) for Airbnb markets in major U.S. cities. This exercise simplifies the Airbnb marketplace in many dimensions, but we believe it is useful for offering insights into the prevalence of estimators with the wrong sign in real-world markets. In our investigation, we find that such estimators are plausible in the Airbnb context. Notably, all markets in the study yielded estimators with the wrong sign when we simulated LR experiments, which would be a typical approach for testing pricing changes in practice. Although this analysis is based on a simplified market model, the results suggest that changes of sign are prevalent enough to be consequential and of concern for practitioners. We conclude in Section 9.
@@ 2 Related Work
The phenomenon we study is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts. More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015). Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014, Fradkin, 2019, Liu et al., 2020), showing its importance in practice. Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020, Candogan et al., 2024), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), rollout experiments Han et al., (2023), and message passing (Shirani and Bayati, 2024).
@@
A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), Roemheld and Rao, (2024), Deng and Le, (2023), Simchi-Levi and Wang, (2023). Some are less related to our work; for example, Simchi-Levi and Wang, (2023) focuses on designing price experiments to mitigate the tail risk of revenue loss from experimentation. More related to our paper, both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects). Roemheld and Rao, (2024) uses an empirical study to suggest that false positives may be inflated due to interference in pricing experiments. More similar to our findings, Deng and Le, (2023) also notes the possibility of the sign of the treatment effect estimator being incorrect; in their setting this is due to an effect they call lead-day bias in markets with expiring goods. In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign. Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper—those where a continuous variable (such as price) is incrementally varied—as marginal policy changes, and develops a framework to study them.
@@
The closest of these works to ours is Munro et al., (2021), which studies general estimands in an economic price equilibrium setting, and decomposes the treatment effect into direct and indirect effects; the former is obtained via “canonical” estimators, while the latter corresponds to our bias, and the authors show how it can be corrected using local price experiments. In that work, the direct and indirect effects are obtained as marginal changes with respect to the treatment assignment probability (inspired by work in Hu et al., (2022)), rather than the price as in our present paper. Further, our emphasis in applying our results is on structural insight regarding when changes of sign may be obtained for objectives such as profits or revenue, while their paper is focused on design and estimation for general estimands in market equilibrium. (An earlier version of Munro et al., (2021) considers treatment changes in a continuous variable, with an emphasis on relating the resulting differential treatment effect to direct and indirect effects.)
@@ 3.1 Market Model and Estimand
The firm is interested in optimizing the performance metric by adjusting the decision variable $x$. In particular, the firm is interested in the change in the target metric, when the decision variable takes on a new “treatment” value, denoted by $x_{1}$, moving away from the status-quo “control” value, denoted by $x_{0}$. This difference, $T(x_{1})-T(x_{0})$, corresponds to the global treatment effect ($GTE$). We define a local version of the $GTE$ as the derivative of the target metric, representing the case in which the treatment value of the decision variable is “close” to the control value:
$GTE(x_{0})$ $:=T^{\prime}(x_{0})\ .$ (1)
@@ 3.2 Experiment: Design, Model, and Estimator
Our work focuses on settings where the function $T$ may be unknown to the firm or platform. Therefore, an experiment is conducted to determine the direction in which the decision variable should be adjusted in order to improve the performance metric (a form of policy improvement). We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction $1-q$ to control. We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated to each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
@@
It is worth noting that in the settings we are considering, randomization can occur on either side of the market, i.e., customers or products (or listings). In a customer-side randomization experiment (CR), a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. On the other hand, in a listing-side randomization experiment (LR), a customer is exposed to a mix of treatment and control products (or listings). As before, we denote by $x_{0}$ the baseline (control) value of the decision variable and by $x_{1}$ the treatment value.
@@
We denote the in-experiment performance metrics for control and treatment goods by the functions $T_{0}(x_{0},x_{1},q)$ and $T_{1}(x_{0},x_{1},q)$ respectively. The function $T_{0}(x_{0},x_{1},q)$ (resp., $T_{1}(x_{0},x_{1},q)$) corresponds to the performance metric in the control (resp., treatment) group when the control and treatment values are given by $x_{0}$ and $x_{1}$, respectively, and a fraction $q$ of units are randomized into treatment. The nature of these functions will become more clear in the context of the applications described below. Notably, this notation abstracts away from the specific experimental design (whether CR or LR). We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
@@
The canonical experimental designs at each side of the market have an associated natural “canonical” local estimator. To define this estimator, consider a sequence of experiments for which the treatment value of the decision variable approaches the control value. Then, we compute the relative difference between the performance metrics appropriately normalized by the treatment and control fractions, respectively, which corresponds to the Horvitz-Thompson estimator, or equivalently in our formal mean-field setting, the difference-in-means estimator. More formally, we consider a sequence of experiments indexed by $n\in\mathbb{Z}_{+}$, in which the treatment value approaches the control value: $\lim_{n\to\infty}x_{1}^{n}=x_{0}$. We define a local version of the canonical estimator as:
$\widehat{GTE}(x_{0},q)$ $:=\lim_{n\to\infty}\frac{\frac{T_{1}(x_{0},x_{1}^{n},q)}{q}-\frac{T_{0}(x_{0},x_{1}^{n},q)}{(1-q)}}{x_{1}^{n}-x_{0}}\ .$ (2)
@@
The assumption states that if the values of control and treatment decision variables are the same, then both the scaled control and treatment performance metrics correspond to the market-wide metric. The assumption will play a key role in our analysis. To further understand it, consider $T_{1}$ with an extended domain $0<q\leq 1$ (the argument for $T_{0}$ is symmetric). Now, suppose $T_{1}$ is linear in its third argument over the relevant domain, that is, for any $0<q\leq 1$ and $10<\alpha<$ 1, $T_{1}(x,x,\alpha q)=\alpha T_{1}(x,x,q)$. If we take $q=1$, then $T_{1}(x,x,\alpha)/\alpha=T_{1}(x,x,1)=T(x)$, where the second equation follows directly by definition. Hence, the assumption is implied by constant (i.e., linear) returns of the performance metric in the fraction randomized. Note that in practice, experimental estimates typically implicitly rely on the assumption of constant returns in order to extrapolate estimates for control and treatment groups to the entire market. For the marketplace models and metrics we study later, this assumption will be satisfied.
@@
$\widehat{GTE}(x_{0},q)$ $=\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{1}(x_{0},x_{0},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$
@@
Finally, we define the estimation bias of this local estimator as: [^3] Note that in statistics, bias is often defined as the estimator minus the estimand. However, we make the preceding choice because (as we will see) the resulting sign of $Bias$ is typically nonnegative, making it easier to obtain practical interpretations of our results.
$Bias(x_{0},q)$ $:=GTE(x_{0})-\widehat{GTE}(x_{0},q)$ (4)
$=T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)\ .$ (5)
@@
$\Box$
The result provides an intuitive characterization of bias. Specifically, $T_{0y}(x_{0},x_{0},q)$ and $T_{1x}(x_{0},x_{0},q)$ correspond to the relative changes in performance for the control and treatment groups, respectively, when the decision variable of the other group changes marginally. The cross-partial derivatives represent the effect on the performance metric across experimental groups, making them an intuitive indicator of interference. As expected, as this interference effect increases, bias correspondingly increases.